Indicate whether or not each of the following stru
Answer #1
Concepts and reason

• The compound with cyclic structure (ring), flat with conjugated double bonds present in the ring is said to be aromatic.

• All cyclic compounds cannot be aromatic; there are non-aromatic cyclic compounds also present.

• The structure should be aromatic or non-aromatic. It can be predicted with the help of the 4n+2{\rm{4n + 2}} rule.

Fundamentals

4n+2rule:4n+2=numberofπelectronsn=integer(n=1,2,3,....)\begin{array}{l}\\{\rm{4n + 2}}\,{\rm{rule:}}\\\\{\rm{4n + 2}}\,{\rm{ = }}\,\,{\rm{number}}\,\,{\rm{of}}\,{\rm{\pi }}\,{\rm{electrons}}\\\\\,\,\,\,\,\,\,{\rm{n}}\,\,{\rm{ = }}\,\,{\rm{integer}}\,\,{\rm{(n}}\,{\rm{ = 1,}}\,{\rm{2,}}\,{\rm{3}}\,{\rm{,}}....{\rm{)}}\\\end{array}

Example:

The above structure has 6 π electrons, and the structure is flat with conjugated double bonds. By substituting (n=1)\left( {{\rm{n}}\,{\rm{ = 1}}} \right) in the 4n+2 formula:

4(1)+2=6{\rm{4}}\left( {\rm{1}} \right){\rm{ + }}\,{\rm{2}}\,\,{\rm{ = 6}}

The above structure obeys the 4n+2{\rm{4n + 2}} rule ( 4(1)+2=6{\rm{4}}\left( {\rm{1}} \right){\rm{ + }}\,{\rm{2}}\,\,{\rm{ = 6}} ). Therefore, it is aromatic.

0.
0.
.
Number of it electrons =10
Number of
electrons =8
4(2) + 2 =10
Obeys 4n+2 rule
4(1) + 2 =6
4(2) + 2 =10
Does not obey

B is aromatic and A is not.

B is not aromatic, as it does not obey the 4n+2{\rm{4n + 2}} rule, and A is aromatic because it obeys the 4n+2{\rm{4n + 2}} rule.

Both A and B are aromatic.

B is not aromatic, as it does not obey the 4n+2{\rm{4n + 2}} rule.

A and B are not aromatic.

A is aromatic because it obeys the 4n+2{\rm{4n + 2}} rule.

A is aromatic and B is not.

A is aromatic because it obeys the 4n+2{\rm{4n + 2}} rule.

Ans:

0.
0.

O
A is aromatic, B is not.
O
B is aromatic, A is not.
O
O
Both A and B are aromatic.
A and B are not aromatic.